OpenStudy (anonymous):

Pre Calc: For what value of k will the graph of 6x + ky = 6 be perpendiculuar to the graph of 2x – 6y = 12? A.1/2 B.)4 C.)2 D.)5

OpenStudy (lgbasallote):

first...solve for y in terms of x in 2x - 6y = 12 do you know how?

OpenStudy (anonymous):

@lgbasallote y=x/3-2 correct?

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now...solve for y in terms of x in 6x + ky = 6

OpenStudy (anonymous):

y=6/k-6x/k

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

right. so that has a slope of 6/k for it to be perpendicular to y = x/3 - 2, the slope should be -3 agree?

OpenStudy (anonymous):

@lgbasallote yessss

OpenStudy (lgbasallote):

wait... i saw wrong

OpenStudy (lgbasallote):

the slope of that equation is -6/k

OpenStudy (lgbasallote):

so now.. equate -6/k to -3 and solve for k \[\huge -\frac 6k = -3\]

OpenStudy (anonymous):

k=2 @lgbasallote thank you <3

OpenStudy (lgbasallote):

welcome